Grade 12th passMagical Mathematics[Interesting Approach]

Show that the lines (x – 5)/7 = (y + 2)/-5 = z/1 and x/1 = y/2 = z/3 are perpendicular to each other.

Harsh Singh Chauhan
5 Years agoGrade 12th pass

1 Answer

SJ
5 Years ago
Welcome to Askiitians

Given lines are:

(x – 5)/7 = (y + 2)/-5 = z/1 and x/1 = y/2 = z/3

The direction ratios of the given lines are 7, -5, 1 and 1, 2, 3, respectively.

We know that,

Two lines with direction ratios a1, b1, c1and a2, b2, c2are perpendicular to each other if a1a2+ b1b2+ c1c2= 0

Therefore, 7(1) + (-5) (2) + 1 (3)

= 7 – 10 + 3

=0

Hence, the given lines are perpendicular to each other.

Thanks